OpenStudy (anonymous):

simplify 99 - 3(4 x 2^3)

OpenStudy (angela210793):

99-12-24=63

OpenStudy (anonymous):

thats not right

OpenStudy (angela210793):

Why not? 99-3*4-3*8=99-12-24=63

OpenStudy (angela210793):

Do you get it?

OpenStudy (anonymous):

the test im taking says its not

OpenStudy (angela210793):

W8....ur right 99-3(32)=99-96=3 O'm so sorry i haven't seen it well Sorry the right answer is 3